Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

coin-Probability

Expert replies
by gmat740 » Tue Jul 28, 2009 3:15 am
When tossed, a certain coin has equal probability of landing on either side. If the coin is tossed 3 times, what is the probability that it will land on the same side each time?

A. 1/8
B. 1/4
C. 1/3
D. 3/8
E. 1/2

OA-B

IMO=>Let coin lands on head side
so, P(head) = 1/2
3 times tossed= (1/2)*(1/2)*(1/2) =1/8

But this is not the answer.

Please correct me.
Last edited by gmat740 on Tue Jul 28, 2009 4:39 am, edited 1 time in total.
Join the discussion
Source: — Problem Solving |

by scoobydooby » Tue Jul 28, 2009 3:44 am
would go for B

there are two possibilites:

HHH or TTT

probability of HHH+probability of TTT=1/8+1/8=1/4
Join the discussion

by viju9162 » Tue Jul 28, 2009 7:39 am
I also got 1/8.. I always miss out what scoobydooby has mentioned.. there are equal chances for both head and tail to occur..

hence 1/8 + 1/8 = 1/4 ..

Thanks for posting the question
"Native of" is used for a individual while "Native to" is used for a large group
Join the discussion

Re: coin-Probability

by ankitns » Wed Jul 29, 2009 7:23 pm
The key here is to realize that we DONT care for 3 tosses...but only 2 tosses. It does not matter what the first toss is...but the second and third need to match it....

so say if the first was heads...probability of second to be heads is 1/2 and third to be head is also 1/2....hence 1/2*1/2 = 1/4.

Cheers.
gmat740 wrote:When tossed, a certain coin has equal probability of landing on either side. If the coin is tossed 3 times, what is the probability that it will land on the same side each time?

A. 1/8
B. 1/4
C. 1/3
D. 3/8
E. 1/2

OA-B

IMO=>Let coin lands on head side
so, P(head) = 1/2
3 times tossed= (1/2)*(1/2)*(1/2) =1/8

But this is not the answer.

Please correct me.
Attempt 1: 710, 92% (Q 42, 63%; V 44, 97%)
Attempt 2: Coming soon!
Join the discussion

by shibal » Thu Jul 30, 2009 9:24 am
Basically it can land head or tail, it asks the prob of landing equally in the 3 attempts, hence it can be 3 times haed or tail

1/8 + 1/8 = 1/4
Join the discussion